At a local convenience store, 2 bags of chips and 7 containers of dip cost $23. However,3 bags of chips and 4 containers of dip cost $15. What is the cost of one bag of chips and one container of dip?
Answer by ewatrrr(24785)

 
Hi
Let x and y represent the cost of the chips and dip respectively
Question states***
2x + 7y = $23
3x + 4y = $15
2x + 7y = $23
-2x -8/3y = $15*-2/3 |multiplying 2nd Eq by -2/3 to eliminate x
(13/3)y = $13
y = $3.00 and x = $1.00 |x = ($15-$12)/3
CHECKING our Answer***
2x + 7y = 23
$2 + $21 = $23